OpenStudy (anonymous):

which equation is the equation of a line that passes through (-10, 3) and is perpendicular to y = x - 7 ? A) y = 5x + 3 B) y = - 1/5x - 7 C) y = - 1/5x + 1 D) y = 1/5x + 5

OpenStudy (anonymous):

well unless you meant to put a 5 in front of the x in the original problem none of those works

OpenStudy (anonymous):

a 5

OpenStudy (anonymous):

ok so then the slope would flip over and become negative to become -1/5 then you would plug everything into point-slope form (y-y1)=m(x-x1) y-3=-1/5(x-(-10)) distribute y-3=-1/5x-2 add three y=-1/5x+1 the answer is c